Systems Staff Meeting Notes 7/6/98

Systems Staff Meeting Notes
July 6, 1998

Present: Brian Burton, Bill Covey, Debra Harris, Winston Harris, Suzy Shaw

Announcements
  1. The Network Coordinator position goes to Personnel this week. Expect to start interviewing for the position by the middle of August.
  2. Jason Monahan should be stopping by today - he will be returning for 6 weeks. He will work in Destry's office - organizing the CD ROMs & doing server documentation. He will also work on hardware as need be. This will also be the case for Scott Olipra, since ordering is expected to be slow right now.
  3. Debra has posted the phone extensions for LAC & Special Collections on the System Homepage. Preservation soon to come.
Ad Hoc
  1. Systems approach to providing service - there have been some comments to the effect that some staff members are 'afraid' to call on Systems for help. They feel as though they are bothering people.
  2. We need to come up with a way to avoid intimidation with our annoyance.
  3. Need to not talk down to staff & try not to impose deadlines unless they are far in advance.
  4. Computer Support people are generally feared & loathed - leave the especially troublesome calls to Bill.
  5. Winston has suggested the possibility of evaluation sheets - to be filled out by a staff member following the resolution of a trouble call.
Liaisons
  1. This past meeting was postponed due to the holiday, the next one will be on Thursday, July 16.
Networking
  1. IP management for switches in Music & Resource Services are not working. CISCO is investigating, they think the problem could be on NERDC's end or could be a malfunction with the boxes.
  2. Xerox printing - has been a disaster. It cannot work under NT without an individual sign-on per work station, which is not acceptable for us.
  3. Printing will now go on a Novell 3.11 server, with IPX on all public clients (for now).
  4. The software writer has assured us that it will work on Novell. The server will be known as Smathers PS (print server) located in the Novell neighborhoods.
  5. The equipment is to be set up in our server room this week.
  6. Preservation - wiring to be completed this week. Feed coming from 2 different places: our server room & the 2nd floor communications closet. They will get their phones at this time as well.
  7. When it's complete, there will be a fiber run from Com Closet 1 to our server room, which will link all the equipment in this building.
Software Projects Status
  1. ILL - Winston has met with David Fuller about the migration of his stuff off SASS. Lending may be migrated by September, he needs to check with David again.
  2. Personnel - Winston & Sharon met with Tom, Evelyn & Grace for a general info session regarding their migration. Should be resolved by the Fall.
  3. LASSIE - The request portion of LASSIE should be functional & operational by the end of the month. The prototype is currently installed on Carolyn's machine. Brian is working on security measures.
  4. Timer Software - Brian spoke with John Ashcraft - he wants to find Win API protocol to determine if a computer is running or not.
DP Coordinator
  1. Trouble calls will be routed to whomever Bill designates while Debra is out.
Calendar
  1. Sharon and Lee out today, July 6.
  2. ALA in Washington, D.C. is June 24 - 30.
  3. Rich will be gone to ALA/vacation June 24-July 15.
  4. Debra out July 7 - 10.
  5. Winston tentatively out Friday, July 10.
  6. Bobbie out July 13 - 17
  7. Brian out July 16 & 17.
  8. Lee out Sept. 17-30.


Back to the Systems Homepage

Last updated March 7, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu